01 December 2023 12:28 Repair work on Tank 3 is well under way and running according to schedule so far. The tank has been split and the liner removed from the Linac trench. This is to allow it to be cleaned at the same time as the main tank reducing repair time. 01 December 2023 08:40 Attention TS1 Methane Moderator Users: a liquid transfer will start in around 5 minutes 30 November 2023 08:40 Attention TS1 Methane Moderator Users: a liquid transfer will start in around 5 minutes 29 November 2023 17:21 BEAM OFF - It has been confirmed that tank 3 window has failed. Due to the recent failures and the decreasing time between failures the decision has been made to remove the tank lid for a deep clean. This procedure is expected to take approximately seven days to complete. Therefore, beam is not expected until next Wednesday. Updates on progress will be issued as the task progresses. 29 November 2023 08:40 Attention TS1 Methane Moderator Users: a liquid transfer will start in around 5 minutes 29 November 2023 08:34 Machine update:- MS beam restored to both target stations at 08:28hrs. 29 November 2023 06:15 We are experiencing problems with a beam intensity monitor. An expert is on site and is trying to remedy the fault. Update at 08:00. 29 November 2023 04:32 We have established base rate beam. We will increase beam rep rate as the vacuum in tank 3 improves. 29 November 2023 00:47 An expert has started the conditioning of Tank 3 window. If this goes well we should see low rep rate beam around 04:00 hrs. Update at 04:00. 28 November 2023 20:57 Tank 3 Vacuum pressure is now at a level where we can start RF conditioning. However we can't lock the area up quite yet because we are having issues with the Ion source. Once we have locked the area and our Injector Engineers commence conditioning we will give an estimate of when MS beam will be restored. 28 November 2023 10:11 Machine update:- The RF tank window was replaced yesterday and the vacuum is continuing to improve to normal operational values. We will next update when the vacuum is at a level for conditioning of the tank to commence. 28 November 2023 08:40 Attention TS1 Methane Moderator Users: a liquid transfer will start in around 5 minutes 27 November 2023 20:25 Beam OFF - Unfortunately the RF window in tank 3 has failed again. Engineers are working on the replacement now. Earliest beam to target is tomorrow evening. Update to follow as work progresses. 27 November 2023 18:58 Beam off while we investigate an issue with mod 3. Update at 20:00 27 November 2023 08:40 Attention TS1 Methane Moderator Users: a liquid transfer will start in around 5 minutes 26 November 2023 21:05 MS beam restored to both target stations. 26 November 2023 18:18 Tank 3 conditioning was uneventful and the tank pressure has recovered well. We have established base rate beam to both target stations. We will increase beam rep rate as beamloss improves and the vacuum recovers further. 26 November 2023 14:17 Tank 3 Vacuum has recovered well this morning and we have a Linac Engineer on site conditioning the tank to RF. We hope to get the tank to operational levels with a good vacuum by late afternoon. If all goes well MS beam should be achieved early this evening. If we have any issues we will give an update. 26 November 2023 08:40 Attention TS1 Methane Moderator Users: a liquid transfer will start in around 5 minutes 25 November 2023 19:53 Specialists have confirmed that Tank 3 window has failed. They are currently in the process of replacing the window and will then start pumping down the tank vacuum. This takes approximately 24 hours so there will be no beam until tomorrow evening. We will update on progress tomorrow afternoon. 25 November 2023 17:46 Linac modulator 3 has tripped and we have had a vacuum rise in the tank. At this stage we think the RF window might have failed again. We have an Injector and Vacuum specialist on their way to site to confirm. We will update at 19:00 Hours. 25 November 2023 08:40 Attention TS1 Methane Moderator Users: a liquid transfer will start in around 5 minutes 24 November 2023 22:36 Beam ON 24 November 2023 21:05 The klixon interlocks have been reset and we are currently running up the main magnet power supply which will take approximately 90 minutes. Beam should be re-established around 22:30. 24 November 2023 19:44 The damaged klixon in the outer synchrotron has been replaced. However, we are now having issues resetting the klixon interlocks to enable us to run the main magnet power supply. Investigations into the cause of this are ongoing. Update @ 21:00. 24 November 2023 18:15 We have accessed the Outer Synchrotron to address an issue with one of the Klixons on SP4 QC magnet. We also have a diagnostics issue and an equipment specialist is coming in to look at this for us. Update at 19:30 Hours 24 November 2023 17:23 The Main Magnet Power Supply has tripped off on magnet over temperature. We are accessing the Inner Synchrotron to check water flows to the magnets. Update at 18:00 Hours. 24 November 2023 10:25 We are having issues with an Extract Kicker power supply. An equipment expert is working on the fault. Beam should be restored soon. 24 November 2023 09:10 Beam on at 09:01 Hours 24 November 2023 08:45 The power supply is now running. We are locking up the Inner Synchrotron and hope to restore beam within the next 30 minutes. 24 November 2023 08:40 Attention TS1 Methane Moderator Users: a liquid transfer will start in around 5 minutes 24 November 2023 08:35 We have made some progress with the failed power supply. Update at 09:30 Hours 24 November 2023 07:32 We have had no success repairing the Synchrotron steering magnet power supply. We are currently consulting an equipment expert for advice. Further update at 08:30 Hours. 24 November 2023 06:46 Beam OFF - We are currently replacing the output converter on R3HD1 power supply following replacement of the the CCU unit. Update at 07:15 24 November 2023 06:00 Beam OFF - A trim quad power supply in the synch has failed and requires replacement. Update at 06:00 23 November 2023 08:40 Attention TS1 Methane Moderator Users: a liquid transfer will start in around 5 minutes 22 November 2023 08:40 Attention TS1 Methane Moderator Users: a liquid transfer will start in around 5 minutes 22 November 2023 04:22 Beam OFF - The water temp protection (Klixon) on SP4 QC magnet had actually failed and required replacement. This requires access to the outer synch to replace the klixon before running the MMPS back up again. Update at 05:00 22 November 2023 02:33 Beam OFF - The MMPS has tripped on an over temp interlock. We need to access the inner synch to adjust the water flow before running the MMPS back up. Update at 03:30 22 November 2023 01:52 Beam OFF - Extract septum 1 PSU has tripped off and requires resetting. This should take approximately 30 minutes 21 November 2023 21:17 MS Beam established to both target stations. 21 November 2023 19:10 Tank 3 beamloss is high but the vacuum pressure is slowly improving. We are increasing rep rate as losses allow and hope to get to MS beam within the next couple of hours. 21 November 2023 17:54 The vacuum level on tank 3 is at 5.0E-06mBar. We will attempt some base rate beam to see if we can recover the vacuum to its normal operational value a little quicker. We will give updates on progress. 21 November 2023 13:49 Tank 3 RF conditioning is underway and going well. However, the vacuum level is still too high to run beam so this will need to recover further. MS beam is still not expected until early evening. Update @ 17:00. 21 November 2023 08:40 Attention TS1 Methane Moderator Users: a liquid transfer will start in around 5 minutes 20 November 2023 18:43 We have confirmed that the RF window has failed on Linac Tank 3. The window is currently being replaced. Unfortunately the Vacuum pump down to operational levels on this tank takes approximately 24 Hours. Realistically there will be no beam before early evening Tuesday 21st November. We will give an update on progress tomorrow afternoon. 20 November 2023 17:35 We have a suspected RF window failure on Linac Tank 3. Our Mechanical and Vacuum Engineers are investigating. If a window failure is confirmed we will give an estimate on how long the repair, vacuum pump down and conditioning will take. 18 November 2023 15:27 BEAM ON 18 November 2023 15:07 BEAM OFF Modulator 4 has tripped off on a water flow - We need to run down the Injector and enter in order to check flow rates. Update at 15:45 17 November 2023 18:42 BEAM OFF - We currently have a vacuum rise in SP7 and are in contact with an expert for advice. Update at 20:00 17 November 2023 16:10 BEAM ON 17 November 2023 15:40 Machine update:- The water leak on Dipole 8 has been repaired. We have locked up up the Outer Synchrotron and are currently running up the MMPS. All being well we